OpenStudy (anonymous):

If a/b + b/a = -1. Then find a^3 + b^3.

OpenStudy (shubhamsrg):

here..we easily see a^2 + b^2 = ab now a^3 + b^3 = (a+b) (a^2+ b^2 -ab) so the ans is 0!!

OpenStudy (anonymous):

wouldnt:\[a^2+b^2 = -ab\]?

OpenStudy (shubhamsrg):

oh yes am sorry!! leme figure out something else..

OpenStudy (shubhamsrg):

well in that case you see a^3 - b^3 = 0 thus a=b now should be easier..

OpenStudy (anonymous):

So a^3 + b^3 = a^ + a^3 = 2a^3. But does that help?

OpenStudy (shubhamsrg):

well i dont guess it can be simplified further..

OpenStudy (shubhamsrg):

no no.wait,,theres something fishy...a=b does not satisfy a/b + b/a = -1 hmmn..

OpenStudy (shubhamsrg):

which means a and b are complex..

OpenStudy (anonymous):

i dont think there are any solutions to that equation. if you try to solve:\[x+\frac{1}{x}=-1\]You end up with:\[x^2+x+1=0\]which only have complex solutions.

OpenStudy (anonymous):

yeah, a and b would have to be complex.

OpenStudy (shubhamsrg):

do this..let a/b =x so you have a quadratic eqn x^2+ x + 1 =0 thus you solve a/b = something,,so,,a = b(something) try and make that substitution..dont know if this'd be a great help//

OpenStudy (anonymous):

in fact, i think:\[\frac{a}{b}=-\frac{1}{2}+i\frac{\sqrt{3}}{2}\]and :\[\frac{b}{a}=-\frac{1}{2}-i\frac{\sqrt{3}}{2}\]works.

OpenStudy (shubhamsrg):

from (a+b)(a^2 + b^2 -ab) ,,factor like this : a(1 + b/a) * (a^2)(1 + (b/a)^2 - b/a ) now try.. hmmn..

OpenStudy (shubhamsrg):

note that b^2 = a and a^2 =b

OpenStudy (anonymous):

yes thats right.

OpenStudy (anonymous):

first we get a^2 +b^2 = -ab now putting this in a^3 +b^3 we get a+b*(-2ab) then using (a+b)^2.... a+b=\[\sqrt{ab}\]...so ans. should be -2[ab\[\sqrt{?}\]]

OpenStudy (anonymous):

sqrt(ab)..*......nt "?"